# Large $N$ limit of Wilson loops on orientable closed surfaces in the light of Koike-Schur-Weyl duality and Spin Networks

### 11 Mar 2026 v1

@@ 1. Introduction
where $\tilde{\ell}$ denotes the lift of a loop $\ell\in\mathrm{L}(\Sigma)$ for the fundamental cover $\tilde{\Sigma}\to\Sigma$ of $\Sigma$ endowed with the volume form identified with the one of $\Sigma$. The main aim of the current text is to prove this conjecture for closed orientable surface of genus $\geq 2.$
@@
Another one builds on recent technical progress [Mag22, Mag25] on the large $N$ limit of $\mathbb{E}[W_{\ell}]$ under the Atiyah-Bott-Goldman measure when $G=SU(N)$. The latter is the weak limit of the Yang-Mills measure as $vol(\Sigma)\to 0$ and can be identified [KS94] as the normalised Liouville volume form for the symplectic form on the moduli space of flat connexions over $\Sigma$ [Gol84, AB83]. The current work’s technical improvement consists in showing convergence of $\mathbb{E}[|W_{\ell}|^{2}]$ under the Atiyah-Bott-Goldman measure and to generalise it the Yang-Mills measure. Following111the first steps of this approach based on weighted surface sums, have also been revisited in [CPS25] for the closely related lattice Yang-Mills measure with Wilson action. [MP19], the strategy of [Mag25] we build on here, is to write explicit formulas for Wilson loops moments as $N^{\chi}$-weighted sums over surfaces with boundaries, with Euler characteristic $\chi$ which can be suitably upper-bounded. As in [MP23] and [Mag25], the latter bound depends crucially on the geometry of the loops, a refinement of Dehn’s algorithm [BS87] and the realisation of $SU(N)$-irreducible representations as traceless tensors.
@@ 2.1. Heat kernel on (special) unitary groups
In this text, $N\geq 1$ is an integer, $G_{N}$ or simply $G$ stands for the compact Lie group given either by $U(N)=\{U\in M_{N}(\mathbb{C}):UU^{*}=\mathrm{Id}_{N}\}$ or by $SU(N)=\{U\in U(N):\det(U)=1\}$. Its Lie algebra $\mathfrak{g}$ is respectively $\mathfrak{u}(N)=\{X\in M_{N}(C):X+X^{*}=0\}$ or $\mathfrak{su}(N)=\{X\in\mathfrak{u}(N):\mathrm{Tr}(X)=0\}.$ It is endowed with its unique Haar probability measure $dg$ and the inner product
@@
$\Delta_{G}=\sum_{i=1}^{\dim(\mathfrak{g})}\mathcal{L}_{X_{i}}^{2}$
@@ 4.3. $\Omega$-functions and combinatorial expression for the expansion of the partition function
For the combinatorial problem, one of the difficulty of these expressions comes from definition and interpretation of the group algebra element $\Omega_{n,m}^{2-2g}$. When $g=1,$ this term disappears and Riemann-Hurwitz formula is enough to interpret the $\frac{1}{N}$-expansion for the Witten-Zeta functions in terms of Hurwitz numbers counting simple ramified covering of the torus. [^20] This is derived with complete arguments for instance in [LM25, Nov24] in the chiral case and also appeared e.g. in [CMR97]. In the chiral case and general $g\geq 1,$ since the $\frac{1}{N}$ expansion of the standard Weingarten element $\Omega^{-1}_{n,0}$ is known [Col03, CS06, Nov10, L0́8], it is also possible to give such expressions considering for instance mixed monotone Hurwitz numbers [Nov24].
@@ 4.4.5. Extension
When $n,m\geq 1,$ for any diagram $\pi\in\mathcal{D}_{\varepsilon}$, consider the diagram $\pi^{n,m}$ obtained by replacing each string with $n+m$ copies, changing signs of the of vertices in the last $m$ copies. More precisely, for all $v\in[s(n+m)],$ writing $v=r+q(n+m)$ for some $q\in[s]$ and $1\leq r\leq n,$ set $\tilde{\varepsilon}(v)=\varepsilon(q)$ when $r\leq n$ and $\tilde{\varepsilon}(v)=-\varepsilon(q)$ if $n<r\leq n+m,$ and for $\eta\in\{-1,1\},$ set
@@ 4.4.6. Second moment diagram for one relator
It is then elementary to check from Lemma 4.9 the following.
@@
$\dim(V^{\lambda})\dim(V^{\mu})|\mathrm{Tr}(\omega(U_{i}))|^{2}\chi_{\lambda,\mu}(r(U_{1},\ldots,U_{k}))$ $=\mathrm{Tr}_{\mathcal{T}_{\varepsilon,N}}(Q^{\lambda,\mu}\pi^{n,m}_{r,\omega}U^{\otimes\mathfrak{a}})$
$=\mathrm{Tr}_{\mathcal{T}_{\varepsilon,N}}(Q^{\lambda,\mu}\pi^{n,m}_{r,\omega}Q^{\lambda,\mu}U^{\otimes\mathfrak{a}}).$
where $\varepsilon:[|r|(n+m)+2|\omega|]\to\{-1,1\},\mathfrak{a}:[|r|(n+m)+2|\omega|]\to\mathcal{A}$ are the juxtaposition of the sign function and arc colouring of $\omega$ and $\omega^{-1}$, with the extension of the ones of $r,$ while
(144) $Q^{\lambda,\mu}=(P_{N}^{\lambda,\mu})^{\otimes r}\otimes\mathrm{id}_{2|\omega|}.$
@@
$\dim(V^{\lambda})\dim(V^{\mu})$ $|\mathrm{Tr}(\omega(U_{i})A)|^{2}\chi_{\lambda,\mu}(r(U_{1},\ldots,U_{k})B)$
@@ 4.4.7. Unitary integration and invariant ideal
When $\varepsilon:[s]\to\{-1,1\}$ and $\varepsilon^{-1}(1)=\#\varepsilon^{-1}(-1),$ we denote by $S_{\varepsilon}$ the set of bijections from $\varepsilon^{-1}(-1)$ to $\varepsilon^{-1}(1).$ When $\mathfrak{a}:[s]\to\mathcal{A}$ is an arc colouring, $S_{\mathfrak{a}}$ denotes the subset of bijections $\alpha\in S_{\varepsilon}$ with
@@
for any $v\in\varepsilon^{-1}(-1).$ The standard unitary Weingarten formula (100) and bounds on the Weingarten function (see e.g. [CS06, Cor. 2.4. and Prop. 2.6.]) reads as follows in our notations. When $\sigma\in S_{n}$ is a permutation denote by $\sigma$ its number of cycles.
@@
We can now deduce our main integral representation formula in terms of Brauer algebra.
@@
Note that horizontal diagrams span an ideal $\mathcal{I}_{\varepsilon}$ of $\mathcal{B}_{\varepsilon}(N).$ When $n=m,$ $\mathcal{I}_{\varepsilon}\not=\{0\}$ and when $2n\geq N$, it follows from Koike-Schur-Weylduality [Hal96] that $\mathcal{B}_{\varepsilon}(N)$ is semi-simple while $\mathcal{I}_{\varepsilon}$, as a sub-algebra, is simple and isomorphic to $\mathrm{End}(V^{[\emptyset,\emptyset]})$ where $V^{[\emptyset,\emptyset]}=\mathrm{Hom}_{U(N)}(V_{\rho_{N}},\mathcal{T}_{n,m}).$
@@ 4.5. Surfaces associated to products of Brauer diagrams
Let us call Brauer map the data of a tuple $\mathfrak{m}=(\tau_{u},\pi,\tau_{b},\varphi)$ formed of a coloured walled Brauer diagram $\pi\in\mathcal{D}_{\varepsilon}$ together with a horizontal $\varphi\in\mathcal{D}^{h}_{\varepsilon}$ and two compatible diagrams $\tau_{u},\tau_{b}\in\mathcal{D}_{\mathfrak{a},\mathfrak{b}},$ where $\mathfrak{a}:[s]\to\mathcal{A}$ and $\mathfrak{b}:[s]\to\mathcal{S}$ are the arc-boundary colourings and $\varepsilon:[s]\to\{-1,1\}$ is the sign function of $\mathfrak{a}.$ For such a triple, consider the $2$-CW complex built in the following way. Informally, the complex is built from the graph appearing in the product $\tau_{u}\pi\tau_{b}\varphi$ in $\mathcal{B}_{\varepsilon}$ (including possible inner loops) by adding vertical edges between top and bottom vertices of $\varphi,$ idenfitying top and bottom vertices of $\tau_{u}\pi\tau_{b}\varphi$ and gluing disc along cycles associated to the trace of $\tau_{u}\pi\tau_{b}\varphi$ and $\varphi$ in $\mathcal{B}_{\varepsilon}.$ See Figure 1 for an example.
@@
Orient each strand of Brauer diagram $\nu\in\mathcal{D}_{\varepsilon}$ from $v$ to $v^{\prime}$ when $v=(s,\eta),v^{\prime}=(s^{\prime},\eta^{\prime})$ with $\varepsilon_{s}\eta=-1,\varepsilon^{\prime}_{s}\eta^{\prime}=1$. It induces an orientation of its cycles. When $\nu$ is an element of a Brauer map $(\tau_{u},\pi,\tau_{b},\varphi)$, the orientation of $\nu$ is the same as the one induced by $\mathcal{C}_{II}$ circles when $\nu=\varphi$ and by $\mathcal{C}_{I}$-circles otherwise.
Remark 4.12.
Each inner point of a $\varphi$-interval belongs to exactly one $\mathcal{C}_{I}$ circle and one $\mathcal{C}_{II}$ circle, so that the image of an open $\varphi$-interval in $X_{\tau_{u},\pi,\tau_{b},\varphi}$ is included in exactly two $2$-cells and belongs to the interior of $\Sigma_{\mathfrak{m}}$. For any other interval, inner points belong to exactly one circle and hence to the boundary of $\Sigma_{\mathfrak{m}}$. Lastly each point of $V^{\prime}$ is exactly in two intervals of type $\iota,\pi$ or $\tau.$ The union of images of such intervals is a disjoint union of simple loops that is exactly the boundary $\partial\Sigma_{\mathfrak{m}}.$
@@
Remark 4.13.
When $\varphi=\pi_{\alpha,\beta},$ connected components of $\mathcal{C}_{II}$ are in bijection with cycles of $\alpha^{-1}\beta$ so that $\varphi$ is the number of connected components of $\mathcal{C}_{II}.$ Moreover, for a given connected component, its number of $\varphi$-intervals is twice the size of the associated cycle of $\alpha^{-1}\beta$. Orienting $\varphi$-intervals as in remark 4.11, top and bottom strands have opposite orientation so that the orientation of $\varphi$-intervals alternates along any $\mathcal{C}_{II}$ circle.
@@
Assume $G_{N}$ is $U(N)$ or $SU(N).$ For $\lambda\vdash n,\mu\vdash m,$ any word $r,\omega$ in $\mathcal{A}$, define $\pi_{r,\omega}^{n,m}$ and its colouring and sign functions $\mathfrak{a},\mathfrak{b},\varepsilon$ as in section 4.4.7. Then there is a constant $K>0$ depending only on $n,m,r,\omega,$ such that for all $N\geq 1,$ $U_{1},\ldots,U_{k}\in G_{N},$
@@
if for any $l\in\{u,b\},$ $\tau_{l}=Z_{u,i}\otimes\ldots\otimes Z_{l,|r|}\otimes\mathrm{id}_{2|\omega|}$ for some $Z_{l}\in\mathcal{D}_{\varepsilon}^{|r|}$ and $\kappa^{\lambda,\mu}_{N}(\tau,\pi_{\alpha,\beta})=0$ otherwise. Putting together Corollary 4.4 and (152),
(156) $\mathbb{E}_{U}[|\mathrm{Tr}(\omega(U_{i})A)|^{2}\chi_{\lambda,\mu}(r(U_{1},\ldots,U_{k})B)]=\sum_{(\tau_{u},\pi,\tau_{b},\varphi)\in\mathcal{M}_{n,m,r,\omega}}\kappa^{\lambda,\mu}_{N}(\tau_{u},\tau_{b},\varphi)N^{\#\mathcal{C}_{I}}\mathrm{tr}_{I}(A,B)$
@@
$\mathrm{tr}_{I}(A,B)=N^{-\#c.c.\,\mathcal{C}_{I}}\mathrm{Tr}_{\mathcal{T}_{\varepsilon,N}}(\tau_{u}\pi\tau_{b}\varphi V^{\otimes\varepsilon})$
and $V_{1},\ldots,V_{|r|+2|\omega|}\in U(N)$ are as in Corollary 4.4. Since $\mathrm{Tr}_{\mathcal{T}_{\varepsilon,N}}(\tau_{u}\pi\tau_{b}\varphi V^{\otimes\varepsilon})$ is a product of $\mathcal{C}_{I}$ traces in words in $V_{1},\ldots,V_{s}$ and their inverse, [^24] More explicitely, for each $\mathcal{C}_{I}$-cycle $c$, consider the unitary matrix $h_{c}\in U(N)$ formed by multiplying resp. by $B,B^{-1},A$ or $A^{-1}$ whenever $c$ visits $n]\times\{1\},\{n+1,\ldots,n+m\}\times\{2\},\{|r|(n+m)+1\}\times\{2\}$ or $\{|r|(n+m)+|\omega|+1\}\times\{2\}.$ Then $\mathrm{tr}_{I}(A,B)=\prod_{c\,\,\mathcal{C}_{I}-\text{cycle}}\frac{1}{N}\mathrm{Tr}(h_{c}).$ $|\mathrm{tr}_{I}(A,B)|\leq 1.$ Now thanks to (4.11) and Lemma 4.2, there is $K>0$ such that for any Brauer map $(\tau_{u},\pi,\tau_{b},\varphi)$ with type I and II cycles $\mathcal{C}_{I}$ and $\mathcal{C}_{II},$
(157) $|\kappa^{\lambda,\mu}_{N}(\tau_{u},\tau_{b},\varphi)|N^{\#c.c.\,\mathcal{C}_{I}}\leq KN^{-s+\#c.c.\,\mathcal{C}_{II}+\#c.c\,\mathcal{C}_{I}}=KN^{\chi(\Sigma_{\tau_{u},\pi,\tau_{b},\varphi})}.$
@@
Assume $\omega$ is a cyclically shortest length representative of an element of $\Gamma_{g}$ and $r$ is the surface word (161). Assume $G_{N}$ is $U(N)$ or $SU(N).$ For any $k\geq 1,$ there is $K_{k}>0$ such that for all $\alpha\in\mathbb{Z}^{\downarrow N}$ with $|\alpha|\leq k,$ and all $N\geq 1,$ $A,B\in G_{N},$

### 31 Mar 2026 v2

@@ 1. Introduction
where $\tilde{\ell}$ denotes the lift of a loop $\ell\in\mathrm{L}(\Sigma)$ for the fundamental cover $\tilde{\Sigma}\to\Sigma$ of $\Sigma$ endowed with the volume form identified with the one of $\Sigma$. The main result of the current text is to prove this conjecture for closed orientable surface of genus $\geq 2.$
@@
Another one builds on recent technical progress [Mag22, Mag25] on the large $N$ limit of $\mathbb{E}[W_{\ell}]$ under the Atiyah-Bott-Goldman measure when $G=SU(N)$. The latter is the weak limit of the Yang-Mills measure as $vol(\Sigma)\to 0$ and can be identified [KS94] as the normalised Liouville volume form for the symplectic form on the moduli space of flat connexions over $\Sigma$ [Gol84, AB83]. A technical improvement of the current work consists in showing convergence of $\mathbb{E}[|W_{\ell}|^{2}]$ under the Atiyah-Bott-Goldman measure and to generalise it the Yang-Mills measure. Following111the first steps of this approach based on weighted surface sums, have also been revisited in [CPS25] for the closely related lattice Yang-Mills measure with Wilson action. [MP19], the strategy of [Mag25] we build on here, is to write explicit formulas for Wilson loops moments as $N^{\chi}$-weighted sums over surfaces with boundaries, with Euler characteristic $\chi$ which can be suitably upper-bounded. As in [MP23] and [Mag25], the latter bound depends crucially on the geometry of the loops, a refinement of Dehn’s algorithm [BS87] and the realisation of $SU(N)$-irreducible representations as traceless tensors.
@@ 2.1. Heat kernel on (special) unitary groups
In this text, $N\geq 1$ is an integer, $G_{N}$ or simply $G$ stands for the compact Lie group given either by $U(N)=\{U\in M_{N}(\mathbb{C}):UU^{*}=\mathrm{Id}_{N}\}$ or by $SU(N)=\{U\in U(N):\det(U)=1\}$. Its Lie algebra $\mathfrak{g}$ is respectively $\mathfrak{u}(N)=\{X\in M_{N}(\mathbb{C}):X+X^{*}=0\}$ or $\mathfrak{su}(N)=\{X\in\mathfrak{u}(N):\mathrm{Tr}(X)=0\}.$ It is endowed with its unique Haar probability measure $dg$ and the inner product
@@
$\Delta_{\mathfrak{g}}=\sum_{i=1}^{\dim(\mathfrak{g})}\mathcal{L}_{X_{i}}^{2}$
@@ 4.3. $\Omega$-functions and combinatorial expression for the expansion of the partition function
For the combinatorial problem, one of the difficulty of these expressions comes from definition and interpretation of the group algebra element $\Omega_{n,m}^{2-2g}$. When $g=1,$ this term disappears and Riemann-Hurwitz formula is enough to interpret the $\frac{1}{N}$-expansion for the Witten-Zeta functions in terms of Hurwitz numbers counting simple ramified covering of the torus. [^20] This is derived with complete arguments in [LM25, Nov24] in the chiral case and in [LM25b] in the non-chiral case, and also appeared e.g. in [CMR97]. In the chiral case and general $g\geq 1,$ since the $\frac{1}{N}$ expansion of the standard Weingarten element $\Omega^{-1}_{n,0}$ is known [Col03, CS06, Nov10, L0́8], it is also possible to give such expressions considering for instance mixed monotone Hurwitz numbers [Nov24].
@@ 4.4.5. Extension
When $n,m\geq 1,$ for any diagram $\pi\in\mathcal{D}_{\varepsilon}$, consider the diagram $\pi^{n,m}$ obtained by replacing each string with $n+m$ copies, changing signs of the vertices in the last $m$ copies. More precisely, for all $v\in[s(n+m)],$ writing $v=r+q(n+m)$ for some $q\in[s]$ and $1\leq r\leq n,$ set $\tilde{\varepsilon}(v)=\varepsilon(q)$ when $r\leq n$ and $\tilde{\varepsilon}(v)=-\varepsilon(q)$ if $n<r\leq n+m,$ and for $\eta\in\{-1,1\},$ set
@@ 4.4.6. Second moment diagram for one relator
The following is elementary to check from Lemma 4.9.
@@
$d^{\lambda}d^{\mu}|\mathrm{Tr}(\omega(U_{i}))|^{2}\chi_{[\lambda,\mu]_{N}}(r(U_{1},\ldots,U_{k}))$ $=\mathrm{Tr}_{\mathcal{T}_{\varepsilon,N}}(Q^{\lambda,\mu}\pi^{n,m}_{r,\omega}U^{\otimes\mathfrak{a}})$
$=\mathrm{Tr}_{\mathcal{T}_{\varepsilon,N}}(Q^{\lambda,\mu}\pi^{n,m}_{r,\omega}Q^{\lambda,\mu}U^{\otimes\mathfrak{a}}),$
where $\varepsilon:[s]\to\{-1,1\},\mathfrak{a}:[s]\to\mathcal{A}$ are the juxtaposition of the sign function and arc colouring of $\omega$ and $\omega^{-1}$, with the extension of the ones of $r,$ while
(144) $Q^{\lambda,\mu}=(P_{N}^{\lambda,\mu})^{\otimes|r|}\otimes\mathrm{id}_{2|\omega|}.$
@@
$d^{\lambda}d^{\mu}$ $|\mathrm{Tr}(\omega(U_{i})A)|^{2}\chi_{[\lambda,\mu]_{N}}(r(U_{1},\ldots,U_{k})B)$
@@ 4.4.7. Unitary integration and invariant ideal
When the sign function satisfies $\varepsilon:[s]\to\{-1,1\}$ and $\varepsilon^{-1}(\{1\})=\#\varepsilon^{-1}(\{-1\}),$ denote by $S_{\varepsilon}$ the set of bijections from $\varepsilon^{-1}(\{-1\})$ to $\varepsilon^{-1}(\{1\}).$ When $\mathfrak{a}:[s]\to\mathcal{A}$ is an arc colouring, $S_{\mathfrak{a}}$ denotes the subset of bijections $\alpha\in S_{\varepsilon}$ with
@@
for any $v\in\varepsilon^{-1}(\{-1\}).$ The standard unitary Weingarten formula (100) and bounds on the Weingarten function (see e.g. [CS06, Cor. 2.4. and Prop. 2.6.]) reads as follows in our notations. When $\sigma\in S_{n}$ is a permutation denote by $\sigma$ its number of cycles.
@@
We can now deduce our main integral representation formula in terms of Brauer algebra using Corollary 4.2.
@@
Note that horizontal diagrams span an ideal $\mathcal{I}_{\varepsilon}$ of $\mathcal{B}_{\varepsilon}(N).$ When $n=m,$ $\mathcal{I}_{\varepsilon}\not=\{0\}$ and when $2n\leq N$, it follows from Koike-Schur-Weyl duality [Hal96] that $\mathcal{B}_{\varepsilon}(N)$ is semi-simple while $\mathcal{I}_{\varepsilon}$, as a sub-algebra, is simple and isomorphic to $\mathrm{End}(V^{[\emptyset,\emptyset]})$ where $V^{[\emptyset,\emptyset]}=\mathrm{Hom}_{U(N)}(V_{\rho_{N}},\mathcal{T}_{n,m}).$
@@ 4.5. Surfaces associated to products of Brauer diagrams
Let us call Brauer map the data of a tuple $\mathfrak{m}=(\tau_{u},\pi,\tau_{b},\varphi)$ formed of a coloured walled Brauer diagram $\pi\in\mathcal{D}_{\varepsilon}$ together with a horizontal $\varphi\in\mathcal{D}^{h}_{\varepsilon}$ and two compatible diagrams $\tau_{u},\tau_{b}\in\mathcal{D}_{\mathfrak{a},\mathfrak{b}},$ where $\mathfrak{a}:[s]\to\mathcal{A}$ and $\mathfrak{b}:[s]\to\mathcal{S}$ are the arc-boundary colourings and $\varepsilon:[s]\to\{-1,1\}$ is the sign function of $\mathfrak{a}.$ For such a tuple, consider the $2$-CW complex built in the following way. Informally, the complex is built from the graph appearing in the product $\tau_{u}\pi\tau_{b}\varphi$ in $\mathcal{B}_{\varepsilon}$ (including possible inner loops) by adding vertical edges between top and bottom vertices of $\varphi,$ idenfitying top and bottom vertices of $\tau_{u}\pi\tau_{b}\varphi$ and gluing discs along cycles associated to the trace of $\tau_{u}\pi\tau_{b}\varphi$ and $\varphi$ in $\mathcal{B}_{\varepsilon}.$ See Figure 1 for an example.
@@
Remark 4.11 (Orientation).
Orient each strand of a Brauer diagram $\nu\in\mathcal{D}_{\varepsilon}$ from $v$ to $v^{\prime}$ when $v=(s,\eta),v^{\prime}=(s^{\prime},\eta^{\prime})$ with $\varepsilon_{s}\eta=-1,\varepsilon^{\prime}_{s}\eta^{\prime}=1$. It induces an orientation of its cycles. When $\nu$ is an element of a Brauer map $(\tau_{u},\pi,\tau_{b},\varphi)$, the orientation of $\nu$ is the same as the one induced by $\mathcal{C}_{II}$-circles when $\nu=\varphi$ and by $\mathcal{C}_{I}$-circles otherwise.
@@
When $\varphi=\pi_{\alpha,\beta},$ connected components of $\mathcal{C}_{II}$ are in bijection with cycles of $\alpha^{-1}\beta$ so that $\varphi$ is the number of connected components of $\mathcal{C}_{II}.$ Moreover, for a given connected component, its number of $\varphi$-intervals is twice the size of the associated cycle of $\alpha^{-1}\beta$. Orienting $\varphi$-intervals as in remark 4.11, top and bottom strands have opposite orientation so that the orientation of $\varphi$-intervals alternates along any $\mathcal{C}_{II}$-circle.
Remark 4.13 (Boundary).
Each inner point of a $\varphi$-interval belongs to exactly one $\mathcal{C}_{I}$-circle and one $\mathcal{C}_{II}$-circle, so that the image of an open $\varphi$-interval in $X_{\tau_{u},\pi,\tau_{b},\varphi}$ is included in exactly two $2$-cells and belongs to the interior of $\Sigma_{\mathfrak{m}}$. For any other interval, inner points belong to exactly one circle and hence to the boundary of $\Sigma_{\mathfrak{m}}$. Lastly each point of $V^{\prime}$ is exactly in two intervals of type $\iota,\pi$ or $\tau.$ The union of images of such intervals is a disjoint union of simple loops that is exactly the boundary $\partial\Sigma_{\mathfrak{m}}.$
@@
Assume $G_{N}$ is $U(N)$ or $SU(N).$ For $\lambda\vdash n,\mu\vdash m$ and any words $r,\omega$ in $\mathcal{A}$, define $\pi_{r,\omega}^{n,m}$ and its colouring and sign functions $\mathfrak{a},\mathfrak{b},\varepsilon$ as in section 4.4.7. Then there is a constant $K>0$ depending only on $n,m,r$ and $\omega,$ such that for all $N\geq 1,$ $U_{1},\ldots,U_{k},A,B\in G_{N},$
@@
if for any $l\in\{u,b\},$ $\tau_{l}=Z_{u,i}\otimes\ldots\otimes Z_{l,|r|}\otimes\mathrm{id}_{2|\omega|}$ for some $Z_{l}\in\mathcal{D}_{\varepsilon}^{|r|}$ and $\kappa^{\lambda,\mu}_{N}(\tau_{u},\tau_{b},\pi_{\alpha,\beta})=0$ otherwise. Putting together Corollary 4.4 and (152),
(156) $\mathbb{E}_{U}[|\mathrm{Tr}(\omega(U_{i})A)|^{2}\chi_{\lambda,\mu}(r(U_{1},\ldots,U_{k})B)]=\sum_{(\tau_{u},\pi,\tau_{b},\varphi)\in\mathcal{M}_{n,m,r,\omega}}\kappa^{\lambda,\mu}_{N}(\tau_{u},\tau_{b},\varphi)N^{\#\text{c.c.}\,\mathcal{C}_{I}}\mathrm{tr}_{I}(A,B)$
@@
$\mathrm{tr}_{I}(A,B)=N^{-\#\text{c.c.}\,\mathcal{C}_{I}}\mathrm{Tr}_{\mathcal{T}_{\varepsilon,N}}(\tau_{u}\pi\tau_{b}\varphi V^{\otimes\varepsilon})$
and $V_{1},\ldots,V_{|r|+2|\omega|}\in U(N)$ are as in Corollary 4.4. Since $\mathrm{Tr}_{\mathcal{T}_{\varepsilon,N}}(\tau_{u}\pi\tau_{b}\varphi V^{\otimes\varepsilon})$ is a product of $\mathcal{C}_{I}$ traces in words in $V_{1},\ldots,V_{s}$ and their inverse, [^24] More explicitely, for each $\mathcal{C}_{I}$-cycle $c$, consider the unitary matrix $h_{c}\in U(N)$ formed by multiplying resp. by $B,B^{-1},A$ or $A^{-1}$ whenever $c$ visits $n]\times\{1\},\{n+1,\ldots,n+m\}\times\{2\},\{|r|(n+m)+1\}\times\{2\}$ or $\{|r|(n+m)+|\omega|+1\}\times\{2\}.$ Then $\mathrm{tr}_{I}(A,B)=\prod_{c\,\,\mathcal{C}_{I}-\text{cycle}}\frac{1}{N}\mathrm{Tr}(h_{c}).$ $|\mathrm{tr}_{I}(A,B)|\leq 1.$ Now thanks to (4.11) and Lemma 4.2, there is $K>0$ such that for any Brauer map $(\tau_{u},\pi,\tau_{b},\varphi)\in\mathcal{M}_{n,m,r,\omega},$
(157) $|\kappa^{\lambda,\mu}_{N}(\tau_{u},\tau_{b},\varphi)|N^{\#c.c.\,\mathcal{C}_{I}}\leq KN^{-s+\#c.c.\,\mathcal{C}_{II}+\#c.c.\,\mathcal{C}_{I}}=KN^{\chi(\Sigma_{\tau_{u},\pi,\tau_{b},\varphi})}.$
@@
Assume $\omega$ is a cyclically shortest length representative of an element of $\Gamma_{g}\setminus\{1\}$ and $r$ is the surface word (161). Assume $G_{N}$ is $U(N)$ or $SU(N).$ For any $k\geq 1,$ there is $K_{k}>0$ such that for all $\alpha\in\mathbb{Z}^{\downarrow N}$ with $|\alpha|\leq k,$ and all $N\geq 1,$ $A,B\in G_{N},$